Which of these physical quantities has base units kgm2s-2?

A. Pressure B. Energy C. Power D. Momentum


1
Expert's answer
2021-10-31T18:11:00-0400
"{\rm[Pressure]=N/m^2=(kg\cdot m/s^2)/m^2}\\
\rm=kg\cdot m^{-1}\cdot s^{-2}"

"{\rm[Energy]=J=N\cdot m=(kg\cdot m/s^2)\cdot m}\\
\rm=kg\cdot m^{2}\cdot s^{-2}"

"{\rm[Power]=J/s=N\cdot m/s=(kg\cdot m/s^2)\cdot m/s}\\
\rm=kg\cdot m^{2}\cdot s^{-3}"

"\rm[Momentum]=kg\cdot m/s=kg\cdot m\cdot s^{-1}"

Answer: B. Energy